MONT 104N – Modeling the Environment Second Writing Assignment Due: No later than 5:00pm on Friday, November 22 November 14, 2019

Background

As you all should be aware from various comments by Prof. Reents at Natural World Cluster events and from me in class, we are planning for our seminar to be participating this coming spring in a cluster-wide Earth Lab Project. The Natural World Cluster faculty are thinking of this as a way to:

• highlight the ongoing climate crisis,

• educate the Holy Cross community about the issues involved,

• continue the work begun by the College administration and staff to address these issues in concrete ways on campus, and

• (possibly) to join in actions by local groups to address these issues in the wider Worcester community (or even more widely).

Nothing quite like this has ever been tried before as a cluster-wide project in the Montserrat program. This means that we are all in unfamiliar territory and we will all need to “buy into” the idea to make this a successful effort(!) It will require a significant commitment of class time, a rethinking of the assignments for the courses in the spring, and certainly a great deal of flexibility and improvisation as the spring semester progresses. But, based on our reading of The Uninhabitable Earth, the Natural World Cluster faculty is convinced that the time is now to try this because of the seriousness of the situation and the need to get people thinking and acting to address it.

We are currently thinking that our first activity in the spring (immediately after we return from the semester break) will involve an “audit” of the use of various resources by the residents of Brooks-Mulledy–things such as:

• water usage for toilets, showers, etc.

• total travel by individuals, teams, etc.

• total energy usage for heating, lighting, electronics, etc.

• total generation of trash and other waste from food packaging, drink containers, waste paper, etc.

• total energy cost of the food consumed by residents based on dietary choices, etc.

The goal there will be simply to get a clearer picture of how we are using resources and what our total carbon footprint is.

When we have gathered and analyzed the results, then we will try to identify whether things can or should be changed, what we can agree to, and what we might want to propose to the campus and the administration. It seems natural to me that, given the topics our class has studied and will study in the spring, one clear role for us would be to lead in analyzing the data collected by the “audit” mentioned above.

As I said above, we will all need to “buy into” this idea to make it work, and for that reason, I am asking you in this writing assignment to reflect on the following questions and to share your ideas for what you want this project to accomplish. It might seem contradictory, but I want you both to “dream big” and to be hard-headed (i.e. realistic).

(1) First and maybe most importantly, are you willing to commit to this “Earth Lab” Natural World Cluster common project idea? Granted, a lot of the specifics of what that will mean are not clear yet. However, it’s important to know this now, because unless you are willing to participate, then there’s not too much point in continuing to plan and develop ideas. This would not be just another meaningless assignment for a class that has no connection with your real life!

If your answer to (1) is “no,” then the assignment is to write an essay of about 3 pages to explain your reaction and to criticize the ideas presented above. If you want to pursue this option, I would suggest that you think of your essay as an attempt to convince others that this is a bad idea by mounting the best arguments you can. That is, don’t just say “this is a stupid idea” – say specifically why you think it is stupid and try to convince the reader that they should think it is stupid too by presenting a well-reasoned argument backed up by appropriate evidence! You can use any evidence you want, but as always, document your sources. The “they say, I say” model would be extremely appropriate for this.

On the other hand, if your answer to (1) is “yes,” your essay (again about 3 pages) should be based around answers to these questions:

(2) Why do you think this is a potentially interesting or valuable thing to do in our class (and with the other seminars in the Natural World Cluster in Montserrat) this coming spring?

(3) Are there other aspects of our environmental impact should we be considering rather than or besides the ones listed above as the targets of our initial “audit”?

(4) If you attended the Sunrise Movement presentation on November 13, what do you think about their model of mass political action by young people to address climate change? Can or should we do something like that or participate in their actions as part of this project?

(5) What sorts of public activities might the Earth Lab project lead to? For instance, there is an annual Academic Conference held at Holy Cross toward the end of the spring semester where results of student research carried out in that academic year are presented. This will occur on April 22 this year. Coincidentally, this is Earth Day, and the date of one of the upcoming Climate Strikes(!) Can we work toward a presentation at the Conference and/or an activity coordinated with that campus event?

General points

You should aim for a lively style and the writing should be in your own voice. Make your essay something you want other people to read. On the other hand, please think twice before using extremely colloquial ways of saying things. If they contribute to the point you are trying to make, that is fine. But do it for a reason. Your sentences should be clear, readable, and not overly wordy. Also, please do not use a thesaurus to find impressivesounding words when simpler alternatives are available. Errors in mechanics (spelling, punctuation, grammar) and usage should not annoy the reader. Try as much as possible to avoid those mistakes.